# Accelerated Sequential Flow Matching: A Bayesian Filtering Perspective

### v1

@@ 5.3 State Estimation
Figure 2: Inference latency v.s. performance for maze planning under varying sampling steps (1, 2, 3, 4, 5, 8, 10, 20 and 50 steps). DF: Pyramid and DF: Full-sequence refer to two denoising schedules of diffusion forcing. Sequential DF uses full sequence denoising.
@@ 6 Conclusion and Limitations
Acknowledgments
We would like to thank Prof. Tailin Wu, as well as Long Wei and Haodong Feng for their insightful discussions and feedback on earlier versions of this work.
@@ Appendix A Deferred Proof and Derivation
(A) independent Gaussian coupling. Suppose $x_{t}(0)\sim p(x_{t}|z_{\leq t})$ and $x_{t}(1)\sim\mathcal{N}(0,I)$ are independently sampled. Let $p_{\text{Gaussian}}$ be the distribution of $\hat{x}_{t}(0)=x_{t}(1)-v(x_{t}(1),1;z_{\leq t})$ using one-step sampling by ODE Eq. 5. Then, $\mathcal{W}^{2}_{2}(p_{\text{Gaussian}},p)=\operatorname{Var}(x_{t}|z_{\leq t}),$ where $\mathcal{W}_{2}(p,q)$ is the 2-Wasserstein distance.
(B) temporally-correlated coupling. Suppose $(x_{t-1},x_{t})\sim p(x_{t-1}|z_{\leq t-1})p(x_{t}|x_{t-1},z_{\leq t})$, and $x_{t}(0)=x_{t},x_{t}(1)=x_{t-1}$. Let $p_{\text{Bayes}}$ be the distribution of one-step sampling by ODE Eq. 6. Then
@@ B.2 Implementation Details of Algorithms 1,2
Forecasting. In forecasting task we have $x_{t}=s_{t+1:t+H}$ and $z_{t}=x_{t}$, where $s_{t}$ is the real physical states we want to predict. To apply Algorithm 1, we take a small finetuning dataset $\{s^{(i)}_{1:T},z^{(i)}_{1:T}\}_{i=1,2,...,n}$ and leverage a pretrained Rectified Flow model $v_{\theta_{0}}$ to generate some trajectories $\hat{s}^{(i)}_{t+1:t+H}\sim p_{\theta_{0}}(s_{t+1:t+H}|z^{(i)}_{\leq t})$. We could directly use $\hat{s}^{(i)}_{t:t+H-1}$ as the source distribution and $s^{(i)}_{t+1:t+H}$ as the target distribution in sequential flow ODE Eq. 6. In practice, to better align the prediction at the same physical time, we instead drop the first $\hat{s}^{(i)}_{t}$ (as it is already observed at time $t$) and pad a $\hat{s}^{(i)}_{t+H-1}$ (can be seen as a moving average) and construct $(\hat{s}^{(i)}_{t+1:t+H-1},\hat{s}^{(i)}_{t+H-1})$ as the source distribution instead.
@@ C.1 State Estimation
C.2 Long-horizon Behavior
Figure 3 compares the RMSE as a function of the forecast lead time h=1,2,,Hh=1,2,...,H. While the auto-regressive model performs better at short horizons, its error increases more rapidly with horizon than flow-based models. In contrast, the sequential flow model with one sampling step performs competitively to the full-step pretrained model across the entire forecast window.
Figure 3: RMSE of Burgers’ equation as a function of forecast lead time.
C.3 Performance-Latency Trade-offs

### v2

@@ 5.4 Performance and Latency Tradeoffs
Figure 3: Inference latency (x axis) v.s. performance (y axis) for maze planning under varying sampling steps (1, 2, 3, 4, 5, 8, 10, 20 and 50 steps). DF: Pyramid and DF: Full-sequence refer to two denoising schedule of diffusion forcing. Sequential DF uses full sequence denoising by default.
@@ 6 Conclusion and Limitations
Acknowledgements
We would like to thank Prof. Tailin Wu, as well as Ruiqi Feng, Long Wei and Haodong Feng for their insightful discussions and feedback on earlier versions of this work.
@@ Appendix A Deferred Proof and Derivation
(A) independent Gaussian coupling. Suppose $x_{t}(0)\sim p(x_{t}|z_{\leq t})$ and $x_{t}(1)\sim\mathcal{N}(0,I)$ are independently sampled. Let $p_{\text{Gaussian}}$ be the distribution of $\hat{x}_{t}(0)=x_{t}(1)-v(x_{t}(1),1;z_{\leq t})$ using one-step sampling by ODE Eq. (5). Then, $\mathcal{W}^{2}_{2}(p_{\text{Gaussian}},p)=\operatorname{Var}(x_{t}|z_{\leq t}),$ where $\mathcal{W}_{2}(p,q)$ is the 2-Wasserstein distance.
(B) temporally-correlated coupling. Suppose $(x_{t-1},x_{t})\sim p(x_{t-1}|z_{\leq t-1})p(x_{t}|x_{t-1},z_{\leq t})$, and $x_{t}(0)=x_{t},x_{t}(1)=x_{t-1}$. Let $p_{\text{Bayes}}$ be the distribution of one-step sampling by ODE Eq. (6). Then
@@ B.2 Implementation Details of Algorithms 1,2
Forecasting. In forecasting task we have $x_{t}=s_{t+1:t+H}$ and $z_{t}=x_{t}$, where $s_{t}$ is the real physical states we want to predict. To apply Algorithm 1, we take a small finetuning dataset $\{s^{(i)}_{1:T},z^{(i)}_{1:T}\}_{i=1,2,...,n}$ and leverage a pretrained Rectified Flow model $v_{\theta_{0}}$ to generate some trajectories $\hat{s}^{(i)}_{t+1:t+H}\sim p_{\theta_{0}}(s_{t+1:t+H}|z^{(i)}_{\leq t})$. We could directly use $\hat{s}^{(i)}_{t:t+H-1}$ as the source distribution and $s^{(i)}_{t+1:t+H}$ as the target distribution in sequential flow ODE Eq. (6). In practice, to better align the prediction at the same physical time, we instead drop the first $\hat{s}^{(i)}_{t}$ (as it is already observed at time $t$) and pad a $\hat{s}^{(i)}_{t+H-1}$ (can be seen as a moving average) and construct $(\hat{s}^{(i)}_{t+1:t+H-1},\hat{s}^{(i)}_{t+H-1})$ as the source distribution instead.
@@ C.1 State Estimation
C.2 Performance-Latency Trade-offs